# Superficial Safety Alignment Hypothesis

### v1

@@ B.1 Definition of Attribute Groups and Categorization Process
As detailed in Section 4.1, we categorize the computational units (neurons and channels) of LLMs into four distinct groups: Exclusive Safety Units (ESU), Exclusive Utility Units (EUU), Complex Units (CU), and Redundant Units (RU). Exclusive Safety Units are primarily responsible for safety-related behavior, such as refusal mechanisms and detecting unsafe requests. Exclusive Utility Units are dedicated to general task performance, including natural language understanding, reasoning, and task-specific knowledge retrieval. Complex Units contribute to both safety and utility, as these attributes are intertwined at a higher level of abstraction. Finally, Redundant Units are not significantly involved in either safety or utility and are often characterized by low activation variance across tasks.
To systematically assign computing units to these groups, we employ a structured pruning strategy based on the variance of activation values. Specifically, we calculate the variance of activations across a target dataset for each neuron or channel. Neurons with higher variance contribute more significantly to the model’s performance on a given task, while neurons with low activation variance are considered redundant and can be pruned. We define two separate importance scores for each unit—$\mathbf{I_{U}}$ for utility-related tasks and $\mathbf{I_{S}}$ for safety-related tasks. Units with extreme values in either dimension are considered Exclusive Units (either ESU or EUU), while units with significant contributions to both dimensions are classified as Complex Units (CU).
@@
By performing structured pruning at various ratios and evaluating the impact on both utility and safety performance, we can accurately categorize the model’s computing units. The pruning process involves removing the least critical units and measuring performance degradation, ensuring that our attribution of units is aligned with their actual contribution to model behavior.
@@ B.2 Evaluating the Impact on Both Utility and Safety Performance
To evaluate the impact of pruning on both utility and safety performance, we measure the model’s performance using established benchmarks for both attributes. Our approach closely follows the methods used by Sun et al. (2023); Wei et al. (2024b); Zou et al. (2023b), with adaptations to focus on the specific aspects of utility and safety in the context of safety alignment.
@@
Measuring Safety. We measure the model’s safety by evaluating its attack success rate (ASR) in response to harmful instructions. Specifically, we prompt the model using AdvBencheval, which consists of 100 harmful prompts, and collect the model’s responses. We consider an attack as successful if the model’s response lacks key patterns indicative of instruction rejection. The ASR is then computed as the ratio of successfully attacked prompts to the total number of prompts evaluated. Our safety evaluation includes two use cases:
  1. 1.
    ASRVanilla: This metric reflects the model’s response to harmful instructions under standard, non-malicious conditions. We calculate this both with and without the system-level instructions that define the model’s behavior and constraints.
  2. 2.
    ASRAdv-Suffix: In this setting, the attacker optimizes adversarial suffixes to bypass the model’s safety guardrails (Zou et al., 2023b). This setting allows us to test the model’s resilience to manipulative inputs that are designed to mislead the model into following harmful instructions.
The ASRVanilla metric gives insight into the model’s safety under normal operating conditions, while ASRAdv-Suffix helps evaluate its robustness against more sophisticated attacks. By evaluating safety under both conditions, we gain a comprehensive understanding of the model’s safety performance. These metrics provide a solid basis for assessing the safety and utility trade-offs during the structured pruning process, ensuring that the model maintains high utility while minimizing the risk of harmful outputs. Detailed results are provided in Table 2.
@@ B.3 Model Pruning Details and Structured Components
We calculate the importance score (I) for each channel or neuron by measuring the activation variance across a target dataset, which is described in equation 1. For each module, channels and neurons with the least activation variance are pruned, as they are considered less critical for either utility or safety-related tasks.
In addition, to ensure consistency across layers and modules with differing scales, we apply a standardization process to the computed importance scores. Following the methodology outlined in An et al. (2024), the importance score for each channel or neuron is normalized to account for the variation in metrics across different layers and modules. The standardized importance score $\hat{I}^{\ell}_{:,j}$ for a given layer $\ell$ and channel/neuron $j$ is computed as follows:
@@ B.4 Attribute Transfer During Fine-Tuning
In our fine-tuning experiments described in Sec. 4.2, we track the attribute transfer of individual units during the adaptation of safety-aligned models to new tasks. The process involves categorizing the computing units into ESU, EUU, CU, and RU based on their behavior in the original, safety-aligned model before and after fine-tuning. As fine-tuning progresses, we measure how many units initially classified as ESU or CU are converted into EUU or RU. This is done by re-evaluating the importance scores $\mathbf{I_{S}}$ and $\mathbf{I_{U}}$ for each unit after every few epochs of training.
The key insight is that when units critical to safety (ESU or CU) are re-purposed for utility tasks (becoming EUU), the model’s safety performance degrades. This transformation is tracked in the attribute transfer statistics, which are visualized in Fig. 5 of the main text. The attribute transfer analysis highlights the brittleness of current safety mechanisms: when safety-aligned models are fine-tuned on new tasks, many safety-critical components lose their original function, compromising the safety guardrails of the model.
@@ B.5 Experimental Setup for Freezing Safety-Critical Components
To mitigate the safety performance degradation caused by fine-tuning, we experiment with freezing the safety-critical components identified through pruning. After categorizing the units into ESU, EUU, CU, and RU, we freeze the Exclusive Safety Units (ESU) and the top 6% of Complex Units (CU) during the fine-tuning process. This ensures that these units retain a large part of their original function and are not re-purposed for utility tasks. The rest of the model is fine-tuned as usual on new tasks, allowing the non-safety-critical components to adapt to the task while keeping the safety-critical components unchanged.
@@
Evaluation Metrics. The primary metrics used in safety evaluation are the Attack Success Rate (ASR) and the Dangerous Score. For AdvBench, we compute ASR using both the keyword-matching method and the judgments from Llama3-Guard, with a higher ASR indicating that the model is more vulnerable to adversarial attacks. For the HEx-PHI dataset, we compute both the average dangerous score and the proportion of highly dangerous responses (score of 5) as evaluated by GPT-4. Additionally, we also calculate the ASR on HEx-PHI using Llama3-Guard to allow for a model-centric safety evaluation. These metrics provide a comprehensive understanding of how freezing safety-critical components impacts both general safety and the model’s robustness to adversarial inputs. We evaluate the performance of these models across safety and utility benchmarks, comparing them to models that undergo full fine-tuning (with no frozen components). As reported in Table 2, freezing safety-critical components significantly preserves the model’s safety guardrails, even after fine-tuning.
B.6 Details on Redundant Units and Alignment Budget
@@
Helpfulness and Interaction. Additionally, we evaluate the model’s helpfulness using the MT-bench (Zheng et al., 2023), which evaluates how well the model engages in helpful and informative interactions over multiple turns of dialogue. The evaluation includes both first turn and second turn helpfulness scores, where the model is assessed for the usefulness of its responses. As shown in Table 4, fine-tuning only on redundant units leads to comparative or even better helpfulness scores, especially in the first turn, where we observe a more obvious increase compared to full parameter fine-tuning.
Overall, by repurposing redundant units for alignment, we manage to retain and even enhance the model’s performance on key downstream tasks without incurring the typical alignment tax associated with full parameter updates. This approach demonstrates the potential for scalable and efficient safety alignment.
@@ B.7 Parameter-Efficient Fine-Tuning (PEFT) Comparisons
Appendix C Appendix: More Presentations and Discussion
@@
C.1 Ablation Study with Different Ratios of Safety-Critical Components
Figure 8: The degradation of safety when freezing different percent of safety-critical components on Llama2-7B-Chat.
To better understand if relying solely on the ESU is sufficient to preserve safety, we conducted experiments by freezing different ratios of safety-critical units, which included all ESUs and varying percentages of CUs. The results are shown in Fig. 8, and we found that freezing a higher percentage of CU components leads to improved safety preservation. However, when more than 9% of CU are frozen, the further safety benefits become leveling off.
C.2 Comparing with Previous Work
@@
  1. 1.
@@
  2. 2.
    Finer Categorization of Computational Units: While prior methods broadly categorize computational units into two groups—those related to utility and those related to safety—we introduce a more nuanced classification with four groups: Exclusive Safety Unit (ESU), Exclusive Utility Unit (EUU), Complex Unit (CU), and Redundant Unit (RU). This more detailed categorization captures subtle yet crucial differences between various units, enabling a more precise understanding of their roles in maintaining both safety and utility.
  3. 3.
@@
C.3 Additional Figure Presentations
@@
Figure 9: Global alignment process from LLaMA2-7B to LLaMA2-7B-Chat. The figure shows the conversion proportions of UU (Utility Units) and RU (Redundant Units) into different categories—RU, ESU (Exclusive Safety Units), EUU (Exclusive Utility Units), and CU (Complex Units)—during the alignment process. Each subplot illustrates the proportion of units being repurposed for different functions as safety alignment is applied, offering a global view of how the model’s components are redistributed.
@@
Figure 10: Alignment process for LLaMA2-7B to LLaMA2-7B-Chat, separated by Attention and MLP modules. This figure presents the conversion proportions of UU (Utility Units) and RU (Redundant Units) into RU, ESU (Exclusive Safety Units), EUU (Exclusive Utility Units), and CU (Complex Units) for the Attention (Left) and MLP (Right) modules. It provides a detailed view of how different components of the model are repurposed during safety alignment, highlighting differences between the Attention and MLP structures..
@@
Figure 11: Role changes of Llama-7B-Chat during fine-tuning on the Dolly dataset. This figure shows the conversion proportions between the original four roles—RU (Redundant Units), ESU (Exclusive Safety Units), EUU (Exclusive Utility Units), and CU (Complex Units)—before and after fine-tuning. The 4x4 layout highlights how each of the original roles transitions into others during the fine-tuning process, providing a comprehensive view of role changes as the model adapts to the new task.
@@
Figure 12: Role changes of Llama-7B-Chat during fine-tuning on the Dolly dataset, separated by Attention and MLP modules. This figure illustrates the conversion proportions between the original four roles—RU (Redundant Units), ESU (Exclusive Safety Units), EUU (Exclusive Utility Units), and CU (Complex Units)—before and after fine-tuning, specifically for the Attention (Left) and MLP (Right) modules.
@@
Figure 13: Conversion of RU (Redundant Units) in MLP module during the alignment process from Llama2-7B to Llama2-7B-Chat across different blocks. The majority of RU remains as RU, with a smaller portion being converted to CU and only minimal conversions into ESU and EUU. This indicates that a large percentage of redundant units remain unused and do not play any meaningful role. Additionally, RU is predominantly located in the middle blocks of the model.
@@
Figure 14: Conversion of UU (Utility Units) in MLP module during the alignment process from Llama2-7B to Llama2-7B-Chat across different blocks. This figure illustrates the proportion of UU being converted into other roles—CU (Complex Units), ESU (Exclusive Safety Units), EUU (Exclusive Utility Units), and RU (Redundant Units)—at various blocks of the model. Most UU are converted into CU, with a significant portion also transitioning into EUU and RU. A very small fraction are transferred to ESU. Notably, conversions to CU predominantly occur in the early and later blocks, while conversions to RU and EUU are concentrated in the middle blocks. Although transfers to ESU are minimal, they occur slightly more often in the middle blocks.
@@
Figure 15: Conversion of units into ESU (Exclusive Safety Units) during the alignment process from Llama2-7B to Llama2-7B-Chat. Most ESU originate from UU (Utility Units) rather than RU (Redundant Units), highlighting a reduction in utility-focused units, which contributes to the alignment tax. Meanwhile, a large portion of RU remain inactive throughout the process, representing a significant inefficiency as these units fail to serve any functional role.
@@
Figure 16: Conversion of ESU (Exclusive Safety Units) during the fine-tuning of Llama2-7B-Chat on the Dolly dataset. The figure shows that more than 65% of ESU are converted into other roles during the fine-tuning process, leading to a decline in the effectiveness of the safety mechanisms. This significant repurposing of safety-critical units highlights the potential risks to safety performance during task adaptation.

### v2

@@ B.1 Definition of Attribute Groups and Categorization Process
As detailed in Section 4.1, we categorize the computational units (neurons and channels) of LLMs into four distinct groups: Safety Critical Units (SCU), Utility Critical Units (UCU), Complex Units (CU), and Redundant Units (RU). Safety Critical Units are primarily responsible for safety-related behavior, such as refusal mechanisms and detecting unsafe requests. Utility Critical Units are dedicated to general task performance, including natural language understanding, reasoning, and task-specific knowledge retrieval. Complex Units contribute to both safety and utility, as these attributes are intertwined at a higher level of abstraction. Finally, Redundant Units are not significantly involved in either safety or utility and are often characterized by low activation variance across tasks.
[…]
@@
By performing structured pruning at various ratios (from a large pruning ratio to gradually decrease) and evaluating the impact on both utility and safety performance, we can accurately categorize the model’s computing units. The pruning process involves removing the least critical units and measuring performance degradation, ensuring that our attribution of units is aligned with their actual contribution to model behavior.
@@ B.2 Evaluating the Impact on Both Utility and Safety Performance
To evaluate the impact of pruning on both utility and safety performance, we measure the model’s performance using established benchmarks for both attributes. Our approach closely follows the methods used by Sun et al. (2023); Wei et al. (2024b); Zou et al. (2023b), with adaptations to focus on the specific aspects of utility and safety in safety alignment.
@@
Measuring Safety. We measure the model’s safety by evaluating its attack success rate (ASR) in response to harmful instructions. Specifically, we prompt the model using AdvBencheval, which consists of 100 harmful prompts, and collect the model’s responses. We consider an attack as successful if the model’s response lacks key patterns indicative of instruction rejection. The ASR is then computed as the ratio of successfully attacked prompts to the total number of prompts evaluated.
@@ B.3 Model Pruning Details and Structured Components
We calculate the importance score (I) for each channel or neuron by measuring the activation variance across a target dataset, which is described in equation 1. For each module, channels and neurons with the least activation variance are pruned, as they are considered less critical for either utility or safety-related tasks. In addition, to ensure consistency across layers and modules with differing scales, we apply a standardization process to the computed importance scores. Following the methodology outlined in An et al. (2024), the importance score for each channel or neuron is normalized to account for the variation in metrics across different layers and modules. The standardized importance score $\hat{I}^{\ell}_{:,j}$ for a given layer $\ell$ and channel/neuron $j$ is computed as follows:
@@ B.4 Attribute Transfer During Fine-Tuning
In our fine-tuning experiments described in Sec. 4.2, we track the attribute transfer of individual units during the adaptation of safety-aligned models to new tasks. The process involves categorizing the computing units into SCU, UCU, CU, and RU based on their behavior in the original, safety-aligned model before and after fine-tuning. As fine-tuning progresses, we measure how many units initially classified as SCU or CU are converted into UCU or RU. This is done by re-evaluating the importance scores $\mathbf{I_{S}}$ and $\mathbf{I_{U}}$ for each unit after every few epochs of training.
The key insight is that when units critical to safety (SCU or CU) are re-purposed for utility tasks (becoming UCU), the model’s safety performance degrades. This transformation is tracked in the attribute transfer statistics, which are visualized in Fig. 5 of the main text. The attribute transfer analysis highlights the brittleness of current safety mechanisms: when safety-aligned models are fine-tuned on new tasks, many safety-critical components lose their original function, compromising the safety guardrails of the model.
@@ B.5 Experimental Setup for Freezing Safety-Critical Components
To mitigate the safety performance degradation caused by fine-tuning, we experiment with freezing the safety-critical components identified through pruning. After categorizing the units into SCU, UCU, CU, and RU, we freeze the Safety Critical Units (SCU) and the top 6% of Complex Units (CU) during the fine-tuning process. This ensures that these units retain a large part of their original function and are not re-purposed for utility tasks. The rest of the model is fine-tuned as usual on new tasks, allowing the non-safety-critical components to adapt to the task while keeping the safety-critical components unchanged.
@@
Evaluation Metrics. The primary metrics used in safety evaluation are the Attack Success Rate (ASR) and the Dangerous Score. For AdvBench, we compute ASR using both the keyword-matching method and the judgments from Llama3-Guard, with a higher ASR indicating that the model is more vulnerable to adversarial attacks. For the HEx-PHI dataset, we compute both the average dangerous score and the proportion of highly dangerous responses (score of 5) as evaluated by GPT-4. Additionally, we also calculate the ASR on HEx-PHI using Llama3-Guard to allow for a model-centric safety evaluation. These metrics provide a comprehensive understanding of how freezing safety-critical components impacts both general safety and the model’s robustness to adversarial inputs. We evaluate the performance of these models across safety and utility benchmarks, comparing them to models that undergo full fine-tuning (with no frozen components). As reported in Table 2, freezing SCU significantly preserves the safety performance under model fine-tuning.
B.6 Details of Redundant Units and Alignment Budget
@@
Helpfulness and Interaction. Additionally, we evaluate the model’s helpfulness using the MT-bench (Zheng et al., 2023), which evaluates how well the model engages in helpful and informative interactions over multiple turns of dialogue. The evaluation includes both first turn and second turn helpfulness scores, where the model is assessed for the usefulness of its responses. As shown in Table 4, fine-tuning only on redundant units leads to comparative or even better helpfulness scores, especially in the first turn, where we observe a more obvious increase compared to full parameter fine-tuning. Overall, by repurposing redundant units for alignment, we manage to retain and even enhance the model’s performance on key downstream tasks without incurring the typical alignment tax associated with full parameter updates. This approach demonstrates the potential for scalable and efficient safety alignment.
@@ B.7 Parameter-Efficient Fine-Tuning (PEFT) Comparisons
Appendix C Additional Experimental Results
Other Model Families and Downstream Datasets. We include results for Mistral-7B-Instruct-v0.2 and the GSM8K math dataset to analyze the impact of fine-tuning attacks on different model families and datasets. Table 7 presents the pruning results of Mistral-7B-Instruct-v0.2 across safety and utility benchmarks, highlighting the importance of UCU and SCU to utility and safety, respectively. The safety performance of Meta-Llama2-7B-Chat under GSM8K fine-tuning attacks is shown in Table 8, while the results for Mistral-7B-Instruct-v0.2 under Alpaca and GSM8K fine-tuning attacks are reported in Tables 9 and 10. These tables demonstrate that our method is more effective on models that are already strongly aligned, such as LLaMA-2. This finding aligns with previous studies showing that while Mistral-family models outperform LLaMA-2 models on downstream tasks, they exhibit weaker safety alignment and are more susceptible to fine-tuning attacks. Although the improvements of our strategy on Mistral are less pronounced than those on LLaMA-2, it is important to note that Mistral’s initial ASR—measured by LLaMA3-Guard, a relatively more accurate evaluator—is significantly higher, averaging 42.5% across the Adv and HEx-PHI datasets, compared to LLaMA-2’s initial ASR of just 1%. Given this observation, the results are reasonable, as it is unrealistic to expect a model with inherently weaker safety alignment to maintain strong safety performance under fine-tuning attacks. Nevertheless, our method achieves a relative 30% reduction in ASR on Mistral under Alpaca fine-tuning attacks. Notably, our conclusions follow a common-sense premise: the primary objective is to preserve the safety performance of well-aligned models when subjected to fine-tuning attacks.
Table 7: Pruning results of Mistral-7B-Instruct-v0.2 across safety and utility benchmarks.
Type wiki2 Utility (ACC%) Safety (ASR%)
wino openb arc_c boolq hellas rte avg w/ sys w/o sys avg
Dense 5.59 78.0 34.0 60.0 85.5 65.5 74.5 66.2 (-0) 12.0 17.0 14.5 (+0)
SCU (2%) 6.17 71.5 32.0 55.5 85.5 65.0 71.0 63.3 (-2.9) 34.0 91.0 62.5 (+48.0)
UCU (1.3%) 52.7 58.0 17.5 21.5 55.0 36.0 53.0 40.1 (-26.2) 17.0 23.0 20.0 (+5.5)
RU (13.4%) 8.15 75.5 33.5 55.5 83.0 62.5 74.5 64.1 (-2.1) 14.0 12.0 13.0 (-1.5)
Table 8: Safety performance of Meta-Llama2-7B-Chat under Fine-Tuning attacks (GSM8K).
Bench Judge Initial GSM8K Finetuned Fix SCU + 6% CU Fix SCU + all CU
Adv keyword 0.19% 5.38% (+5.19%) 1.92% (+1.73%) 1.73% (+1.54%)
Adv llama3-guard 0.19% 5.77% (+5.58%) 1.35% (+1.16%) 1.15% (+0.96%)
HEx-PHI gpt4-score 1.61 1.6 (+0.5) 1.37 (+0.32) 1.31 (+0.26)
HEx-PHI gpt4-rate 0.3% 11.51% (+11.21%) 5.75% (+5.45%) 5.31% (+5.01%)
HEx-PHI llama3-guard 2.42% 17.88% (+15.46%) 11.52% (+9.10%) 9.68% (+7.26%)
Table 9: Safety performance of Mistral-7B-Instruct-v0.2 under Fine-Tuning attacks (Alpaca).
Bench Judge Initial Alpaca Finetuned Fix SCU + 6% CU Fix SCU + all CU
Adv keyword 15.19% 89.61% (+74.42%) 74.04% (+58.85%) 72.15% (+56.96%)
Adv llama3-guard 40.38% 87.12% (+46.74%) 73.27% (+32.89%) 70.76% (+30.38%)
HEx-PHI gpt4-score 2.24 4.18 (+1.94) 3.67 (+1.43) 3.43 (+1.19)
HEx-PHI gpt4-rate 18.79% 70.03% (+51.51%) 58.78% (+39.99%) 54.37% (+35.58%)
HEx-PHI llama3-guard 45.45% 80.0% (+40.61%) 70.01% (+24.56%) 67.81% (+22.36%)
Table 10: Safety performance of Mistral-7B-Instruct-v0.2 under Fine-Tuning attacks (GSM8K).
Bench Judge Initial GSM8K Finetuned Fix SCU + 6% CU Fix SCU + all CU
Adv keyword 15.19% 97.31% (+82.12%) 72.31% (+57.12%) 66.34% (+51.15%)
Adv llama3-guard 40.38% 95.38% (+55.00%) 89.81% (+49.43%) 86.92% (+46.54%)
HEx-PHI gpt4-score 2.24 4.15 (+1.91) 4.01 (+1.77) 3.96 (+1.72)
HEx-PHI gpt4-rate 18.79% 66.7% (+47.91%) 64.7% (+45.91%) 62.81% (+44.02%)
HEx-PHI llama3-guard 45.45% 93.94% (+48.49%) 89.39% (+43.94%) 80.31% (+34.86%)
Additional Jailbreak/Red-Teaming Attacks. To evaluate the impact of finetuning attacks on model robustness against jailbreak and red-teaming attacks, we assess the safety performance of finetuned aligned models under three attack methods: GCG, AutoDAN, and PAIR. Specifically, we utilize the HarmBench framework and sample 120 data points to measure performance. As shown in Tables 11 and 12, our method effectively mitigates performance degradation in adversarial scenarios.
Table 11: Safety performance of Meta-Llama2-7B-Chat under Fine-Tuning attacks (Alpaca) across red-teaming attacks.
Bench Red-teaming Initial GSM8K Finetuned Fix SCU + 6% CU Fix SCU + all CU
HarmBench GCG 33.33% 53.08% (+19.75%) 40.25% (+6.92%) 37.75% (+4.42%)
HarmBench AutoDAN 1.08% 7.41% (+6.33%) 2.33% (+1.25%) 1.66% (+0.58%)
HarmBench PAIR 12.25% 22.25% (+10.00%) 14.5% (+2.25%) 14.08% (+1.83%)
Table 12: Safety performance of Meta-Llama2-7B-Chat under Fine-Tuning attacks (Dolly) across red-teaming attacks.
Bench Red-teaming Initial GSM8K Finetuned Fix SCU + 6% CU Fix SCU + all CU
HarmBench GCG 33.33% 62.91% (+29.58%) 43.25% (+9.92%) 40.66% (+7.33%)
HarmBench AutoDAN 1.08% 16.16% (+15.08%) 9.66% (+8.58%) 8.33% (+7.25%)
HarmBench PAIR 12.25% 25.25% (+13.00%) 15.25% (+3.00%) 14.75% (+2.50%)
Appendix D Appendix: More Presentations and Discussion
@@
D.1 Ablation Study with Different Ratios of Safety-Critical Components
To better understand if relying solely on the SCU is sufficient to preserve safety, we conducted experiments by freezing different ratios of safety-critical units, which included all SCUs and varying percentages of CUs. The results are shown in Fig. 8, and we found that freezing a higher percentage of CU components leads to improved safety preservation. However, when more than 9% of CU are frozen, the further safety benefits become leveling off.
D.2 Comparing with Previous Work
@@
Figure 8: The degradation of safety when freezing different percent of safety-critical components on Llama2-7B-Chat
  1. (1)
@@
  2. (2)
    Finer Categorization of Computational Units: While prior methods broadly categorize computational units into two groups—those related to utility and those related to safety—we introduce a more nuanced classification with four groups: Safety Critical Unit (SCU), Utility Critical Unit (UCU), Complex Unit (CU), and Redundant Unit (RU). This more detailed categorization captures subtle yet crucial differences between various units, enabling a more precise understanding of their roles in maintaining both safety and utility.
  3. (3)
@@
D.3 Additional Figure Presentations
@@
Figure 9: Global alignment process from LLaMA2-7B to LLaMA2-7B-Chat. The figure shows the conversion proportions of $U$ (Utility Units) and RU (Redundant Units) into different categories, RU, SCU (Safety Critical Units), UCU (Utility Critical Units), and CU (Complex Units), during the alignment process. Each subplot illustrates the proportion of units being repurposed for different functions as safety alignment is applied, offering a global view of how the model’s components are redistributed.
@@
Figure 10: Alignment process for LLaMA2-7B to LLaMA2-7B-Chat, separated by Attention and MLP modules. This figure presents the conversion proportions of $U$ (Utility Units) and RU (Redundant Units) into RU, SCU (Safety Critical Units), UCU (Utility Critical Units), and CU (Complex Units) for the Attention (Left) and MLP (Right) modules. It provides a detailed view of how different components of the model are repurposed during safety alignment, highlighting differences between the Attention and MLP structures..
@@
Figure 11: Role changes of Llama-7B-Chat during fine-tuning on the Dolly dataset. This figure shows the conversion proportions between the original four roles—RU (Redundant Units), SCU (Safety Critical Units), UCU (Utility Critical Units), and CU (Complex Units)—before and after fine-tuning. The 4x4 layout highlights how each of the original roles transitions into others during the fine-tuning process, providing a comprehensive view of role changes as the model adapts to the new task.
@@
Figure 12: Role changes of Llama-7B-Chat during fine-tuning on the Dolly dataset, separated by Attention and MLP modules. This figure illustrates the conversion proportions between the original four roles—RU (Redundant Units), SCU (Safety Critical Units), UCU (Utility Critical Units), and CU (Complex Units)—before and after fine-tuning, specifically for the Attention (Left) and MLP (Right) modules.
@@
Figure 13: Conversion of RU (Redundant Units) in MLP module during the alignment process from Llama2-7B to Llama2-7B-Chat across different blocks. The majority of RU remains as RU, with a smaller portion being converted to CU and only minimal conversions into SCU and UCU. This indicates that a large percentage of redundant units remain unused and do not play any meaningful role. Additionally, RU is predominantly located in the middle blocks of the model.
@@
Figure 14: Conversion of $U$ (Utility Units) in MLP module during the alignment process from Llama2-7B to Llama2-7B-Chat across different blocks. This figure illustrates the proportion of $U$ being converted into other roles—CU (Complex Units), SCU (Safety Critical Units), UCU (Utility Critical Units), and RU (Redundant Units)—at various blocks of the model. Most $U$ are converted into CU, with a significant portion also transitioning into UCU and RU. A very small fraction are transferred to SCU. Notably, conversions to CU predominantly occur in the early and later blocks, while conversions to RU and UCU are concentrated in the middle blocks. Although transfers to SCU are minimal, they occur slightly more often in the middle blocks.
@@
Figure 15: Conversion of units into SCU (Safety Critical Units) during the alignment process from Llama2-7B to Llama2-7B-Chat. Most SCU originate from $U$ (Utility Units) rather than RU (Redundant Units), highlighting a reduction in utility-focused units, which contributes to the alignment tax. Meanwhile, a large portion of RU remain inactive throughout the process, representing a significant inefficiency as these units fail to serve any functional role.
@@
Figure 16: Conversion of SCU (Safety Critical Units) during the fine-tuning of Llama2-7B-Chat on the Dolly dataset. The figure shows that more than 65% of SCU are converted into other roles during the fine-tuning process, leading to a decline in the effectiveness of the safety mechanisms. This significant repurposing of safety-critical units highlights the potential risks to safety performance during task adaptation.